take three different phones . find their length and width . write the fraction to their width to length . compare these fractions.
1
Expert's answer
2019-06-21T12:22:34-0400

Samsung Galaxy S10: 149.90 by 70.36 mm. "\frac{width}{length}=\frac{70.36}{149.90}=0.4694."

LG G8 ThinQ: 151.83 by 71.88 mm. "\frac{width}{length}=\frac{71.88}{151.83}=0.4732."

OnePlus 3T: 152.70 by 74.70 mm. "\frac{width}{length}=\frac{74.70}{152.70}=0.4892."

OnePlus 3T has the highest fraction width/length, which is equal to 0.4892.

Samsung Galaxy S10 has the lowest fraction width/length, which is equal to 0.4694.